# Investigating Use Cases of AI-Powered Scene Description Applications for Blind and Low Vision People

### v1

@@ 3. Methods
1
24
Female
Low Vision
N/A
Read with right eye. See colors
Yes
VoiceOver, magnification, and contrast
2
67
Male
Blind
No
Light perception. In extreme conditions identify color.
Yes
VoiceOver
3
57
Male
Blind
Yes
Read with magnification with left eye only.
Yes
CCTV to magnify print, talking thermometer and scale, SeeingAI, and VoiceOver
4
34
Female
Blind
Yes
Quite near-sighted. Sensitivity to light. Able to read with magnification 300%.
Yes
CCTV to magnify print, magnification, VoiceOver, and JAWS
5
55
Male
Blind
No
Only light perception on the periphery
Yes
JAWS,NVDA, VoiceOver, and Voice assistants
6
52
Female
Blind
Yes
No Vision
Yes
VoiceOver, and Voice assistants
7
39
Female
Low Vision
No
Light, color and shape perception. Varies a lot
Yes
ZoomText, and CCTV to magnify
8
44
Male
Blind
Yes
No Vision
Yes
VoiceOver
9
37
Female
Blind
No
No Vision
Yes
JAWS, VoiceOver, and Braille Display
10
39
Male
Blind
Yes
Minimal light perception.
Yes
VoiceOver
11
36
Female
Blind
Yes
No Vision
Yes
VoiceOver, JAWS, and Braille Display
12
32
Female
Low Vision
No
Can see large objects, and colors contrast. Can not read
Yes
JAWS, and VoiceOver
13
39
Female
Blind
No
No Vision
Yes
JAWS, and VoiceOver
14
67
Female
Low Vision
No
Read with magnification.
Yes
JAWS, Braille Display, and magnification
15
47
Female
Blind
No
Light perception, color perception, and shapes perception.
Yes
VoiceOver, and JAWS
16
59
Female
Blind
No
No Vision
Yes
VoiceOver, Voice Assistants, and Braille Display
3.1. Participants
@@ 3.4. Qualitative Analysis
Describe Scene
Gathering visual information about related subjects that form a scene
“Is there a mountain in the background?”
@@
Used in educational or spiritual sites
Schools, churches, and
@@
This photo represents an example of an entry coded for location “Used in healthcare sites”. Participant was located in an optometrist office. A photo of a group of medical equipments for general health checks and an examination table.
Used in recreational or leisure areas
Restaurants, parks, and historical sites
@@
This photo represents an example of an entry coded for location “Used in recreational or leisure areas”. Participant was located in a historical site. A photo of a woman taking a selfie in an open field with an aztec pyramid in the background.
Used in healthcare places
Hospitals, dentist offices, and ophthalmologist offices
@@
This photo represents an example of an entry coded for location “Used in educational or spiritual sites”. Participant was located in a chapel. A photo of an empty chapel with wooden pews.
Used in travel services sites
Bus stops, train stations, and airports
@@
This photo represents an example of an entry coded for location “Used in travel service sites”. Participant was located in an airport. A photo of a gateway waiting area at an airport full of people.
@@
Used in retail businesses
Supermarkets, hair salons, and other businesses
@@
This photo represents an example of an entry coded for location “Used in retail businesses”. Participant was located in a supermarket, at the alcohol section. Shelves full of wine bottles at a supermarket
Used at participant’s work
Any spaces participants defined as their “work”, such as corporate or home offices
@@
This photo represents an example of an entry coded for location “Used at participant’s work”. Participant was in their home studioin recording their podcast. A photo of a black and silve microphone on a stand with several electronic devices on the table
Used in living spaces
Apartments, senior homes, and houses
@@
This photo represents an example of an entry coded for location “Used in living spaces or domiciles”. Participant was located at an assisted living facility. A photo of a stripe patterned drinking cup on a table. In the background there are people on wheel chairs.
Used in unidentified locations
Applied when neither the picture nor the participant’s diary responses could tell us where they were located
@@
This photo represents an example of an entry coded for location “Used in unidentified areas”. Participant was located in a indoor space setting but in their dairy entry they did not mention where they were. A photo of a payphone and a trash can.

### v2

@@ 3.1. Participants
1 24 Female Low Vision N/A Read with right eye. See colors Yes VoiceOver, magnification, and contrast
2 67 Male Blind No Light perception. In extreme conditions identify color. Yes VoiceOver
3 57 Male Blind Yes Read with magnification with left eye only. Yes CCTV to magnify print, talking thermometer and scale, SeeingAI, and VoiceOver
4 34 Female Blind Yes Quite near-sighted. Sensitivity to light. Able to read with magnification 300%. Yes CCTV to magnify print, magnification, VoiceOver, and JAWS
5 55 Male Blind No Only light perception on the periphery Yes JAWS,NVDA, VoiceOver, and Voice assistants
6 52 Female Blind Yes No Vision Yes VoiceOver, and Voice assistants
7 39 Female Low Vision No Light, color and shape perception. Varies a lot Yes ZoomText, and CCTV to magnify
8 44 Male Blind Yes No Vision Yes VoiceOver
9 37 Female Blind No No Vision Yes JAWS, VoiceOver, and Braille Display
10 39 Male Blind Yes Minimal light perception. Yes VoiceOver
11 36 Female Blind Yes No Vision Yes VoiceOver, JAWS, and Braille Display
12 32 Female Low Vision No Can see large objects, and colors contrast. Can not read Yes JAWS, and VoiceOver
13 39 Female Blind No No Vision Yes JAWS, and VoiceOver
14 67 Female Low Vision No Read with magnification. Yes JAWS, Braille Display, and magnification
15 47 Female Blind No Light perception, color perception, and shapes perception. Yes VoiceOver, and JAWS
16 59 Female Blind No No Vision Yes VoiceOver, Voice Assistants, and Braille Display
@@ 3.4. Qualitative Analysis
Describe Scene Gathering visual information about related subjects that form a scene “Is there a mountain in the background?”
@@
Used in educational or spiritual sites Schools, churches, and